OpenStudy (anonymous):

If the ratio of H.M between two positive numbers "a" and "b" (a>b) is to their G.M as 12 to 13 prove that a:b is 9:4

OpenStudy (anonymous):

@ganeshie8 @DLS

OpenStudy (anonymous):

\[\frac{ \frac{ 2ab }{ a+b } }{ \sqrt(ab) }=\frac{ 9 }{4 }\]

OpenStudy (anonymous):

It is equal to 12/13

OpenStudy (anonymous):

2ab/(a+b)/sqrt(ab) =12/13 or 2sqrt(ab) / (a+b) = 12/13

OpenStudy (anonymous):

o yes sorry

OpenStudy (anonymous):

after that

OpenStudy (anonymous):

or (a+b)/2sqrt(ab) = 13/12 thus ((a+b) +2sqrt(ab) ) / ((a+b) +2sqrt(ab) ) =(13+12)/(13-12) or ( sqrt(a) +sqrt(b) )/( sqrt(a) +sqrt(b) ) = 5 or ( sqrt(a) +sqrt(b) ) + ( sqrt(a) - sqrt(b) ) / ( sqrt(a) +sqrt(b) ) - ( sqrt(a) - sqrt(b) ) = (5+1)/(5-1) sqrt(a/b) =6/4 or a/b =9/4

OpenStudy (anonymous):

\[\frac{ 2ab }{ a+b }=\frac{ 12 }{ 13 }\sqrt{ab}\]\[\frac{ 2\sqrt(ab) }{a+b }=\frac{ 12 }{ 13 }\]\[\frac{ \sqrt{ab} }{a+b }=\frac{ 6 }{ 13 }\]

OpenStudy (anonymous):

(( sqrt(a) +sqrt(b) ) + ( sqrt(a) - sqrt(b) ) )/ (( sqrt(a) +sqrt(b) ) - ( sqrt(a) - sqrt(b) ) )

OpenStudy (anonymous):

how can you write 2ab * sqrt(ab) as 2sqrt(ab)

OpenStudy (anonymous):

\[\frac{ a+b }{ \sqrt(ab) }=\frac{ 13 }{ 6 }\]\[\frac{ a }{ \sqrt{ab} }+\frac{ b }{ \sqrt{ab}}\]=13/6

OpenStudy (anonymous):

\[\sqrt{\frac{ a }{ b }}+\sqrt{\frac{ b }{ a }}=\frac{ 13 }{ 6 }\]

OpenStudy (anonymous):

now let \[t=\sqrt{\frac{ a }{ b }}\]\[t+\frac{ 1 }{ t }=\frac{ 13 }{ 6 }\]

OpenStudy (anonymous):

solve for t!!

OpenStudy (anonymous):

t= 3 or t= 2/3

OpenStudy (anonymous):

t=3/2 and t=2/3

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

\[\sqrt{\frac{ a }{ b }}=3/2\]

OpenStudy (anonymous):

a:b=9:4

OpenStudy (anonymous):

b:a cannot be 9/4 due to the constraint given that a>b

OpenStudy (anonymous):

I have some questions i would tell you in two minutees

OpenStudy (anonymous):

I mean questions regarding this question

OpenStudy (anonymous):

@No.name 2ab * sqrt(ab) as 2sqrt(ab) its 2ab/(a+b) / sqrt(ab)/1 = 2ab/ (sqrt(ab)*(a+b)) =2sqrt(ab)/(a+b)

OpenStudy (anonymous):

and @No.name i think matricked has applied compenendo and dividendo

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):